OpenStudy (anonymous):

Ernest has a batting average of 0.340. Write this decimal as a fraction in simplest form. a.17/50 b. 3/10 c. 8/25 d. 9/25

OpenStudy (anonymous):

i am not getting this concept

OpenStudy (turingtest):

340/1000=170/500=85/250=17/50

OpenStudy (turingtest):

look at my other post to your last question. follow it closely.

OpenStudy (anonymous):

ok thanks tuningtest

OpenStudy (turingtest):

I'll post it slightly more slowly 0.2=2/10 0.657=657/100 0.04=4/100 now it can be reduced like a regular fraction

OpenStudy (anonymous):

describe a strategy for finding equivalent fraction